OpenStudy (anonymous):

Need some Algebra 2 help!

OpenStudy (anonymous):

what u need help with sexy

OpenStudy (anonymous):

im doing Alg 2-2 right now too

OpenStudy (anonymous):

Find the solution to the equations. 2x - y = -3 x + y = 0

OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 5) and (-2, 3)

OpenStudy (anonymous):

2x-3y+13=0

rishavraj (rishavraj):

it is 3y-2x-13=0

OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-1, -3) and (2, 1)

OpenStudy (anonymous):

(1 - (-3) / (2 - (-1) (1 + 3) / (2 + 1) 4/3 is your slope y = mx + b (2,1)..x = 2 and y = 1 slope(m) = 4/3 now we sub 1 = 4/3(2) + b 1 = 8/3 + b 1 - 8/3 = b 3/3 - 8/3 = b - 5/3 = b y = 4/3x - 5/3 -- multiply by 3 3y = 4x - 5 -4x + 3y = -5 4x - 3y = 5 4x - 3y - 5 = 0

OpenStudy (anonymous):

balls

OpenStudy (anonymous):

what brought that on ?

OpenStudy (anonymous):

ball is life

OpenStudy (anonymous):

I dont think those are right

OpenStudy (anonymous):

the last one!